A machine gun fires 50gm bullets at a speed of 1000m/s. The gunner, holding the machine gun in his hands, can exert an average force of 180N against the gun. Determine the maximum number of bullets he can fire per minute


1
Expert's answer
2021-06-30T16:19:10-0400

The impulsive force

"I=mv \\

I= 50 \times 10^{-3} \times 1000 = 50"

If man can hold the gun at a recoiling force of 180 N in 1 min

t= 60 s

F= 180 N

Number of bullets per min "= \frac{Ft}{I}"

"= \frac{180 \times 60}{50}=216"

Answer: 216